Synthesize each compound from acetylene. You may use any other organic or inorganic reagents. a.

_________________________________________________________________________________________________________________________________

d. Prepare octane from 1-pentyne, indicate each needed reagents.

e. Synthesize cis-2-hexene from 1-pentyne and an alkyl halide.

f. Synthesize trans-2-hexene from 1-pentyne and an alkyl halide.
Answer #1

Substitution a H-C=C-H 1 mol NaNH 1 H-C=EN -NaB 1 mol NaNH Substitution Br No -NaBI 6 H-C=c-H 1 mol NaNH, VaNH Substitution ~Substitution d = 1 mol NaNH; Na Br -NaBI H.Pd Reduction octane 1 mol NaNH 8 Substitution H3C-B1 -NaBr H, Lindlars catalyst c